# Imaginary Gauge-steerable Edge Modes In Non-Hermitian Aubry-André-Harper Model

### v1

@@ IV Coexistence of boundary-pinned and gauge-steerable edge states
Figure 3: Spectral and real-space structure at the parameter set indicated by the blue star marker in Fig. 2. (a) Open-boundary spectrum in the plane $(\mathrm{Re}\,E,|E|)$, with color representing the Lyapunov exponent. Zoomed windows Z1, Z2, and Z3 mark the bulk state BS.1 and the in-gap states ES.2 and ES.3. In (b) and (c) eigenstates are ordered by increasing $\mathrm{Re}\,E$. (b) Three-dimensional plot of the Hermitian AAH densities $|\phi_{n}|^{2}$ versus eigenstate index and site index. (c) Corresponding plot of the ENHSE–AAH densities $|\psi_{n}|^{2}$ obtained from the same eigenvalues via $\psi_{n}=e^{X_{n}}\phi_{n}$. In (b) and (c) the states BS.1, ES.2, and ES.3 are highlighted. Each eigenstate is normalized to unit norm.
@@
With the nonunitary gauge transformation shown above, we can analyze boundary physics using the Hermitian AAH spectrum as a reference: the open-boundary eigenvalues are unchanged, and for typical zero-mean gauge realizations the spatial decay/growth exponents are inherited from the Hermitian problem. In the ENHSE regime $\lambda<2J$, bulk eigenstates in the continuum are extended in the Hermitian counterpart, yet exponentially decaying solutions can still occur at eigenvalues lying inside bulk gaps AubryAndre1980 , which can be seen in Fig. 3. The gauge transformation retains the corresponding in-gap eigenvalues while reshaping their real-space profiles through the envelope $e^{X_{n}}$. Figure 3(a) shows the OBC spectrum of $\hat{H}$, from which we select one representative bulk state (BS.1) and two representative in-gap modes (ES.2 and ES.3), whose Hermitian counterparts are edge states for the comparison below.
@@
Figures 3(b) and 3(c) compare the real-space distribution of all eigenstates before and after the nonunitary transformation. Fig. 3(b) shows the Hermitian AAH amplitudes $|\phi_{n}|^{2}$ as a function of eigenstate index and site index, in which BS.1 is extended along the chain, while ES.2 and ES.3 are peaked near the left boundary. Fig. 3(c) shows the corresponding ENHSE–AAH amplitudes $|\psi_{n}|^{2}$, obtained from the same eigenvalues via $\psi_{n}=e^{X_{n}}\phi_{n}$ for one realization of the Bernoulli sequence $\{s_{n}\}$. The envelope $e^{X_{n}}$ leaves all eigenvalues and Lyapunov exponents unchanged but strongly reshapes the spatial profiles. For the bulk state BS.1, the transformation produces the familiar ENHSE collective pinning: the Hermitian counterpart state is extended in Fig. 3(b), while its weight in Fig. 3(c) is concentrated near the global maximum of $X_{n}$. The two in-gap modes ES.2 and ES.3 respond in a qualitatively different way. ES.3 remains localized at the edge before and after the transformation: its main peak is always pinned to the left boundary and only its tail is slightly distorted by the envelope. In contrast, ES.2 in Fig. 3(c) is pulled away from the left edge and re-centered at an interior position within the bulk, where $X_{n}$ develops a local maximum. Thus the transformation of ES.2 separates the two roles: the decay exponent remains the intrinsic one set by the Hermitian in-gap edge mode, while the localization center is reselected by the gauge-induced envelope $e^{X_{n}}$ through the local structure of $X_{n}$.
@@
Figure 4: Finite-size diagnostics of the three reference eigenstates in the Hermitian counterpart at $J=\lambda=1$, $\beta=(\sqrt{5}-1)/2$, and $\varphi=\pi/5$. (a) Eigenenergies in the gap hosting ES.2 as a function of the system size NN, with the tracked ES.2 branch highlighted. (b) Same as (a) but for the gap hosting ES.3. (c) Boundary weight $P_{\mathrm{edge}}$ of BS.1, ES.2, and ES.3 versus NN evaluated with a fixed boundary window length $n_{b}=100$. (d) Boundary weight $P_{\mathrm{edge}}$ of BS.1, ES.2, and ES.3 versus the boundary window size $n_{b}$ at $N=40000$.
@@
To establish that the Hermitian modes corresponding to ES.2 and ES.3 are bona fide edge states, i.e., spectrally isolated and localized near the boundaries rather than belonging to the bulk continuum—we perform finite-size diagnostics for the Hermitian AAH chain. We start from a reference chain of size $N=2000$ and increase $N$ up to $20000$ in steps of $\Delta N=500$, keeping $(J,\lambda,\beta,\varphi)$ fixed. We monitor how the in-gap eigenvalues within the windows Z2 (ES.2) and Z3 (ES.3) in Fig. 3(a) evolve with the system size $N$. Figures 4(a) and 4(b) show that the tracked eigenvalues of ES.2 and ES.3 remain inside bulk gaps and do not merge into the neighboring continua as $N$ increases, indicating spectral isolation in the thermodynamic limit. To quantify boundary localization of an eigenstate $\phi_{n}$, we compute the boundary weight
@@
where $n_{b}$ is the window length measured from each boundary. As shown in Fig. 4(c), at fixed $n_{b}=100$ the boundary weight $P_{\mathrm{edge}}$ of the bulk reference state BS.1 decreases as the system size $N$ increases. By contrast, the $P_{\mathrm{edge}}$ values of ES.2 and ES.3 are independent of $N$, consistent with boundary localization. This interpretation is further supported by Fig. 4(d): at fixed $N=40000$, the boundary weights of ES.2 and ES.3 rapidly approach unity as $n_{b}$ increases, whereas BS.1 remains far from saturation. Together, these finite-size diagnostics establish that the Hermitian counterparts of ES.2 and ES.3 are bona fide edge states in the thermodynamic limit. The nonunitary gauge transformation then guarantees that the corresponding OBC eigenvalues are unchanged, while the associated right/left eigenfunctions are reshaped by the gauge-induced envelope, which can shift the probability maximum of ES.2 without altering its in-gap eigenvalue.
The distinct responses of ES.2 and ES.3 to the gauge envelope are robust against changes in the ENHSE realization. To test robustness against the ENHSE realization, we keep $(J,\lambda,\beta,\varphi)$ fixed and vary only the Bernoulli sequence $\{s_{n}\}$. Figure 5 shows representative densities $|\psi_{n}|^{2}$ of ES.2 for five independent realizations. All curves exhibit comparable exponential tails, yet their maxima shift over the chain, tracking prominent local maxima of the cumulative gauge profile $X_{n}$. In contrast, the companion edge mode ES.3 remains essentially pinned to the boundary across realizations. Accordingly, we refer to ES.2 as a gauge-steerable in-gap mode: its localization center can be repositioned by changing the imaginary-gauge configuration, whereas ES.3 behaves as a conventional boundary pinned edge mode that is largely insensitive to the gauge randomness.
Figure 5: Real-space probability densities $|\psi_{n}|^{2}$ of the in-gap state ES.2 for five independent realizations of the Bernoulli sequence $\{s_{n}\}$. All microscopic parameters are identical to those used for ES.2 in Fig. 3, only the random sequence $\{s_{n}\}$ is changed. Each curve corresponds to one realization.
@@ V Local-gain protocol for preparing the gauge-steerable in-gap state
We now show how to selectively prepare the gauge-steerable in-gap mode ES.2 from a bulk initial wave packet by adding weak on-site gain at a single site. The selection mechanism is set by the biorthogonal response of an isolated eigenmode to a local perturbation: for weak gain, the mode with the largest positive imaginary energy shift becomes dominant in the long-time normalized dynamics. Below we show that, for the present model, the relevant biorthogonal weight can be evaluated from the Hermitian reference problem, which fixes the gain site independently of the erratic-gauge realization.
@@
where $\delta v$ is the complex perturbation strength. For an isolated, nondegenerate eigenvalue $E_{\alpha}$, first-order perturbation theory gives Bai2025_SMT_SHE
@@
For a perturbation applied at site $n$, $\chi_{\alpha}(n)$ controls the first-order eigenvalue shift via Eq. (23). We take $\delta v=i\gamma$ with $\gamma>0$ (on-site gain), so that $\mathrm{Im}\,\delta E_{\alpha}=\gamma\,\chi_{\alpha}(n)$ for an isolated mode and $\gamma\ll J$. For a given target mode $\alpha$, we therefore choose
@@
which maximizes its gain-induced imaginary shift and makes it dominant in the long-time normalized dynamics.
@@
with $S$ diagonal in the site basis. It follows immediately that
@@
Therefore the optimal gain site is fixed by the profile of the Hermitian mode and is independent of the erratic-gauge realization.
We now implement the gain-selection rule for the gauge-steerable in-gap state ES.2. Let $|\phi_{\mathrm{ES.2}}\rangle$ denote the corresponding Hermitian eigenstate (with eigenvalue $E_{\mathrm{ES.2}}$) identified in Sec. IV. From Eq. (28) we obtain
@@
The resulting $\chi_{\mathrm{ES.2}}(n)$ is shown in Fig. 6(a), where it attains its maximum at $n=18$; according to Eq. (25), we therefore set the gain site to $n_{0}=18$. With $n_{0}$ fixed, we introduce weak local gain and define
@@
The gain term breaks the exact nonunitary gauge transformation in Sec. III; however, for $\gamma\ll J$ it produces only a small perturbation to the eigenstates, while shifting the eigenvalues according to the biorthogonal weight. In particular, the targeted mode acquires the largest positive imaginary shift,
@@
With our choice of $n_{0}$, ES.2 receives a parametrically larger positive imaginary shift than the remaining modes, as shown in Fig. 6(b), and therefore dominates the long-time normalized dynamics.
Figure 6: (a) Biorthogonal weight $\chi_{\mathrm{ES.2}}(n)$; the star marks the gain site $n_{0}=18$. (b) Complex spectra of $\hat{H}$ (blue) and $\hat{H}_{\mathrm{eff}}$ (red), with a zoom into the gap containing ES.2. (c) Space–time map of the normalized density under $\hat{H}_{\mathrm{eff}}$ for a representative disorder realization, starting from a bulk-centered Gaussian wave packet. (d) Fidelity with respect to ES.2 versus time for five disorder realizations.
@@
We numerically evolve Eq. (32); to remove the trivial overall amplification/decay inherent to non-Hermitian dynamics, the state is normalized to unit norm during the time evolution. As the initial condition we use a normalized Gaussian wave packet centered in the bulk,
@@
with $\sigma=20$. To verify that the dynamics prepares ES.2, we monitor the fidelity with respect to the ES.2 right eigenstate,
$F(t)=\big|\langle\psi^{\mathrm{R}}_{\mathrm{ES.2}}|\Psi(t)\rangle\big|^{2},$ (34)
where $|\psi^{\mathrm{R}}_{\mathrm{ES.2}}\rangle$ is the normalized ES.2 right eigenstate of Hamiltonian (5).
Figure 6(c) shows the evolution for realization 2 of the Bernoulli sequence $\{s_{n}\}$ (the same labeling as in Fig. 5). Starting from the bulk, the normalized profile gradually transfers to the left boundary region and converges to a sharply localized profile whose maximum coincides with the ENHSE-steered peak position of ES.2 for that realization. The fidelity in Fig. 6(d) increases from $F(0)\ll 1$ and saturates close to unity, indicating that the weak local gain preferentially amplifies the ES.2 component. We repeat the same protocol for the five independent ENHSE realizations used in Fig. 5. For each realization, the gain site is kept fixed at the same $n_{0}$ determined once from the Hermitian counterpart, while the long-time peak position follows the realization-dependent envelope $e^{X_{n}}$. The fidelities in Fig. 6(d) all saturate near unity, demonstrating that the preparation protocol is robust to ENHSE randomness and yields a dynamically selected, position-tunable in-gap mode.

### v2

@@ IV Coexistence of boundary-pinned and gauge-steerable edge states
Figure 3: Spectrum and structures for the eigenstates at the parameter set indicated by the blue star marker, i.e., $(J,\lambda)=(1,1)$, in Fig. 2. (a) Spectrum of $(\mathrm{Re}\,E,|E|)$, with color representing the Lyapunov exponent (12) in open boundary condition. Zoomed windows Z1, Z2, and Z3 mark the bulk state BS.1 and the in-gap states ES.2 and ES.3. In (b) and (c), eigenstates are ordered by increasing $\mathrm{Re}\,E$. (b) Distribution of the eigenstates for the Hermitian AAH Hamiltonian (3) versus the eigenstate index $\alpha$ and the site index $n$. (c) Corresponding distribution of the eigenstates for the ENHSE-AAH Hamiltonian (5) with the same eigenvalues. In (b) and (c) the states BS.1, ES.2, and ES.3 are highlighted. Each eigenstate is normalized.
@@
With the nonunitary gauge transformation shown above, we can analyze boundary physics using the Hermitian AAH spectrum as a reference: the eigen-spectrum is maintained in OBC, and for $\bar{h}_{n}=0$ the spatial decay/growth exponents are inherited from the Hermitian counterpart. In the ENHSE regime $\lambda<2J$, bulk eigenstates are extended in the Hermitian counterpart, yet there still exist exponentially decaying eigenstates with eigenvalues lying inside the bulk gaps AubryAndre1980 , which can be seen in Fig. 3. The gauge transformation retains the corresponding in-gap eigenvalues while reshaping the envelopes of the corresponding eigenstates through $e^{X_{n}}$. Figure 3(a) shows the OBC spectrum of $\hat{H}$, from which we select one representative bulk state (BS.1) and two representative in-gap modes (ES.2 and ES.3), whose Hermitian counterparts are edge states for the comparison below.
@@
Figs. 3(b) and 3(c) compare the real-space distribution of all eigenstates before and after the nonunitary transformation in Eq. (5). As shown in Fig. 3(b), the amplitudes $|\phi_{n}|^{2}$ of BS.1 is extended along the chain, while those of ES.2 and ES.3 are peaked near the left boundary. Fig. 3(c) shows the corresponding ENHSE-AAH amplitudes $|\psi_{n}|^{2}$, obtained corresponding to the same eigenvalues via $\psi_{n}=e^{X_{n}}\phi_{n}$ for one Bernoulli sequence. The envelope $e^{X_{n}}$ leaves all eigenvalues and Lyapunov exponents unchanged but reshapes the spatial profiles. For the bulk state BS.1, the transformation produces the familiar ENHSE collective pinning: the state of the Hermitian counterpart behaves extendedly in Fig. 3(b), while its distribution in Fig. 3(c) is concentrated near the global maximum of $X_{n}$. The two in-gap modes ES.2 and ES.3 respond in a qualitatively different manner. ES.3 remains localized at the edge before and after the transformation, i.e., its main peak is always pinned to the left boundary and only its tail is slightly distorted by the envelope. In contrast, ES.2 in Fig. 3(c) is pulled away from the left edge and re-centered at an interior position within the bulk, where $X_{n}$ develops a local peak.
@@
Figure 4: Finite-size adjudication of the three eigenstates BS.1, ES.2 and ES.3 for the Hermitian counterpart when $J=\lambda=1$, $\beta=(\sqrt{5}-1)/2$, and $\varphi=\pi/5$. (a) Eigenenergies in the gap hosting ES.2 versus the system size NN. (b) Same as (a) but for the gap hosting ES.3. (c) Boundary weight $P_{\mathrm{edge}}$ defined in Eq. (19) of BS.1, ES.2, and ES.3 versus NN evaluated with a fixed boundary window length $n_{b}=100$. (d) Boundary weight $P_{\mathrm{edge}}$ of BS.1, ES.2, and ES.3 versus the boundary window size $n_{b}$ at $N=40000$.
@@
To establish that the Hermitian modes corresponding to ES.2 and ES.3 are bona fide edge states, i.e., spectrally isolated and localized near the boundaries rather than merging into the bulk continua with the size increasing, we perform finite-size adjudication for the Hermitian AAH chain. We start from a chain of size $N=1000$ and increase $N$ up to $20000$ in steps of $\Delta N=500$, keeping $(J,\lambda,\beta,\varphi)$ fixed. We monitor how the in-gap eigenvalues within the regions Z2 (ES.2) and Z3 (ES.3) in Fig. 3(a) evolve versus the system size $N$ in Figs. 4(a) and 4(b). It can be seen that the eigenvalues of ES.2 and ES.3 remain inside bulk gaps and do not merge into the neighboring continua as $N$ increases, indicating spectral isolation in the thermodynamic limit.
To quantify the boundary localization of an eigenstate with amplitude $\phi_{n}$ at site $n$, we compute the boundary weight
@@
where $n_{b}$ is the window length measured from each boundary. The larger $P_{\mathrm{edge}}$ means that the eigenstate is more concentrated at the edges of the chain at fixed $n_{b}$. As shown in Fig. 4(c) when $n_{b}=100$, the boundary weight $P_{\mathrm{edge}}$ of the bulk state BS.1 decreases as the system size $N$ increases. By contrast, the $P_{\mathrm{edge}}$ values of ES.2 and ES.3 are independent of $N$. This means ES.2 and ES.3 are boundary-localized, but BS.1 is not. This is further supported by Fig. 4(d): at fixed $N=40000$, the boundary weights of ES.2 and ES.3 rapidly approach unity as $n_{b}$ increases, whereas BS.1 remains far from saturation. Together, these finite-size adjudications establish that the Hermitian counterparts of ES.2 and ES.3 are bona fide edge states. The nonunitary gauge transformation in Eq.(5) then guarantees that the corresponding OBC eigenvalues are unchanged, while the associated right/left eigenfunctions are reshaped by $X_{n}$, which can shift the probability peak of ES.2 without altering its in-gap eigenvalue.
To check the distinct responses of ES.2 and ES.3 to different $X_{n}$, we keep $(J,\lambda,\beta,\varphi)$ fixed and vary only the Bernoulli sequence $\{s_{n}\}$. Fig. 5 shows representative densities $|\psi_{n}|^{2}$ of ES.2 and ES.3 for five independent $\{s_{n}\}$. For ES.2, all curves exhibit comparable exponential tails, yet their peaks shift over the chain, tracking prominent peaks of the cumulative gauge profile $X_{n}$. In contrast, the edge mode ES.3 remains essentially pinned to the boundary. Accordingly, we refer to ES.2 as a gauge-steerable in-gap mode, i.e., its localization center can be repositioned by changing the imaginary-gauge configuration $A(x)$, whereas ES.3 pinned tenaciously at the edge, insensitive to the gauge randomness.
Figure 5: Distributions of the eigenstates ES.2 and ES.3 for five different Bernoulli sequences $\{s_{n}\}$. All parameters are identical to those used for ES.2 and ES.3 in Fig. 3, only the Bernoulli sequence $\{s_{n}\}$ varies. Colored dashed curves represent $|\psi_{n}|^{2}$ of ES.2, black solid curves represent that of ES.3; the corresponding $X_{n}$ are plotted as colored solid curves with semi-transparent filled areas.
@@ V Local-gain protocol for preparing the gauge-steerable in-gap state
To go beyond a static eigenstate analysis, we aim to prepare the gauge-steerable in-gap mode ES.2 from a bulk wave packet using a weak local-gain protocol. Without this gain, the state retains substantial overlap with many bulk modes, so ES.2 is not preferentially selected. We show that weak on-site gain at a single site provides a mode selector: for an isolated eigenmode it induces a small imaginary-energy shift, and the mode with the largest positive shift dominates the long-time normalized dynamics.
@@
where $\delta v$ is a complex perturbation. For an isolated, nondegenerate eigenvalue $E_{\alpha}$, first-order perturbation theory gives Bai2025_SMT_SHE
@@
For such a perturbation, $\chi_{\alpha}(n)$ controls the first-order eigenvalue shift via Eq. (23). We take $\delta v=i\gamma$, i.e., on-site gain, with $J\gg\gamma>0$, so that $\mathrm{Im}\,\delta E_{\alpha}=\gamma\,\chi_{\alpha}(n)$ for an isolated mode. Thus, for a given target mode $\alpha$, we choose
@@
which yields the maximal imaginary shift and this makes the target mode dominant in the long-time normalized dynamics.
@@
with $S$ being diagonal in the site basis. It follows immediately that
@@
Therefore the optimal gain site is fixed by the profile of the Hermitian mode and is independent of the erratic-gauge field.
Based on the analysis above, we now implement the local gain to prepare the gauge-steerable in-gap mode ES.2. Let $|\phi_{\mathrm{ES.2}}\rangle$ denote the corresponding Hermitian eigenstate (with eigenvalue $E_{\mathrm{ES.2}}$) identified in Sec. IV. From Eq. (28) we obtain
@@
The resulting $\chi_{\mathrm{ES.2}}(n)$ is shown in Fig. 6(a), where it attains its maximum at $n=18$; according to Eq. (25), $n_{0}=n$. With $n_{0}$ fixed, we introduce weak local gain and define
@@
The gain term breaks the exact nonunitary gauge transformation in Eq. 5. As $\gamma\ll J$, it produces only a small perturbation to the eigenstates, while shifting the eigenvalues according to the biorthogonal weight. In particular, the targeted mode acquires the largest positive imaginary shift,
@@
With this choice of $n_{0}$, ES.2 receives a parametrically larger positive imaginary shift than the remaining modes, as shown in Fig. 6(b), and therefore dominates the long-time normalized dynamics.
Figure 6: Local-gain protocol for dynamically preparing the gauge-steerable in-gap mode ES.2. (a) Biorthogonal weight $\chi(n)$ of ES.2 in Eq. (29); the star marks the selected gain site $n_{0}=18$. (b) Complex spectra for the unperturbed Hamiltonian $\hat{H}$ defined in Eq. (5)(blue dots) and the gain-perturbed Hamiltonian $\hat{H}_{\mathrm{eff}}$ defined in Eq. (30) (red open circles) for $\gamma=0.05$; the inset zooms into the gap hosting ES.2. (c) Fidelity $F(t)$ to ES.2 for the Bernoulli sequence $\{s_{n}\}$-2 (same labeling as in Fig. 5), for $\gamma=0$ and $\gamma=0.05$. Inset (c1): the evolution of the densities $|\Psi_{n}(t)|^{2}$ for $\gamma=0$ and $\gamma=0.05$. (d) $F(t)$ for five Bernoulli sequences $\{s_{n}\}$-$j$ ($j=1,\dots,5$) used in Fig. 5, evaluated when $\gamma=0.05$ with the same fixed gain site $n_{0}$.
@@
We numerically solve Eq. (32) with normalization of the state during the evolution. A normalized Gaussian wave packet centered in the bulk is chosen as the initial state,
@@
with $\sigma=20$. To confirm the preparation of ES.2, we monitor the fidelity with the right mode of ES.2, $|\psi^{\mathrm{R}}_{\mathrm{ES.2}}\rangle$, versus time,
$F(t)=\big|\langle\psi^{\mathrm{R}}_{\mathrm{ES.2}}|\Psi(t)\rangle\big|^{2}.$ (34)
Figure 6(c) summarizes the dynamics for the Bernoulli sequence $\{s_{n}\}$-2 (the same labeling as in Fig. 5). In Fig. 6(c1), the two space-time density maps compare $\gamma=0$ and $\gamma=0.05$. For $\gamma=0$ (no local gain), the wave packet spreads and retains substantial bulk components rather than converging to ES.2. This is quantified in Fig. 6(c): the fidelity with ES.2 stays significantly below unity throughout the evolution. By contrast, when a weak local gain is applied ($\gamma=0.05$), Fig. 6(c1) shows that the normalized profile is gradually drawn from the bulk toward the left region and eventually converges to a sharply localized pattern whose maximum coincides with the ENHSE-steered peak position of ES.2. Consistently, the fidelity in Fig. 6(c) increases from $F(0)\ll 1$ and saturates to unity, indicating that the local gain selectively enhances the ES.2 component.
To test the generality, we repeat the same protocol for the five independent Bernoulli sequences used in Fig. 5. The gain site is kept fixed at the same $n_{0}$ determined once from the Hermitian counterpart, while the long-time peak position follows the envelope $e^{X_{n}}$. As shown in Fig. 6(d), all fidelities saturate near unity, demonstrating that the preparation works consistently across different erratic-gauge fields and yields a dynamically selected, position-tunable in-gap mode.